OpenStudy (anonymous):

int(x/(4+2*x^2) dx what is the answer? (1/4)*ln(2+x^2) or 1/4 *ln(4+2x^2) ?

OpenStudy (zarkon):

both ( if you add your constant)

OpenStudy (anonymous):

on text the second is the answer whereas on maple the second is answer can you please explain it ?

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

according to the logic the second should be the answer, also when I tried int x/(2*x^2+7*a^2) dx on maple this was the answer (1/4)*ln(2*x^2+7*a^2)

OpenStudy (zarkon):

that is because they are both correct ... \[\frac{1}{4} \ln(4+2x^2)=\frac{1}{4} \ln(2(2+x^2))=\frac{1}{4} \ln(2)+\frac{1}{4} \ln(2+x^2)\] thus the two answers differ by only a constant

OpenStudy (anonymous):

so this difference is trivial?

OpenStudy (anonymous):

i mean that can we ignore it? but why?

OpenStudy (zarkon):

\[\int\frac{x}{4+2x^2}dx\] \[=\frac{1}{4} \ln(4+2x^2)+c_1\] or \[=\frac{1}{4} \ln(2+x^2)+c_2\]

OpenStudy (anonymous):

are both answers same?

OpenStudy (zarkon):

yes...once you include the constant of integration

OpenStudy (anonymous):

if I am not wrong have you abosorbed ln(2) in c2?

OpenStudy (zarkon):

\[\frac{1}{4}\ln(2)\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

got ittttttttttt

OpenStudy (anonymous):

thanks a tonne

OpenStudy (zarkon):

yw